# MINDSTORES: Memory-Informed Neural Decision Synthesis for Task-Oriented Reinforcement in Embodied Systems

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated enhanced capabilities in reasoning (Plaat et al., 2024; Huang & Chang, 2023), planning (Sel et al., 2025), and decision-making (Huang et al., 2024) through methods that strengthen analytical depth. Among the numerous domains of active innovation, the success of AI agents serve as a critical benchmark for assessing our progress toward generally capable artificial intelligence (Brown et al., 2020).
@@
A promising paradigm for such agents leverages LLMs as high-level planners (Jeurissen et al., 2024): the LLM decomposes abstract goals into step-by-step plans (e.g., “mine wood → craft tools → smelt iron”), while a low-level controller translates these plans into environment-specific actions (e.g., movement, object interaction). This “brain and body” architecture capitalizes on the LLM’s capacity for structured reasoning while grounding its outputs in the dynamics of the physical world—a critical capability for real-world applications like robotic manipulation (Shentu et al., 2024; Bhat et al., 2024; Wang et al., 2024b), autonomous navigation (Zawalski et al., 2024), and adaptive disaster response.
InputReview ExperiencePlan GenerationPredict Outcome ExecuteRecord
Failure ManagementNext Iteration
Task: Mine IronReview: Previous experiences for mining needs a pickaxe!Plan: Mine Refer to caption, craft Refer to caption, mine Refer to caption, craft Refer to caption, mine Refer to captionPredict: Probable failure, might die from hunger while finding iron ore.Review: Find food before searching for iron ore.Plan: Mine Refer to caption, craft Refer to caption, mine Refer to caption, craft Refer to caption, craft Refer to caption, hunt Refer to caption, mine Refer to captionPredict: Probable success!
@@
Figure 1: Overview of the MINDSTORES planning architecture. The left shows the iterative experiential learning pipeline leveraging the experience database. Database-related methods are in orange, planning steps are in green, and Minecraft steps are in red. The right shows an example applies this pipeline to an example task in Minecraft.
@@
Related approaches like Voyager (Wang et al., 2023) also fall short of experiential learning. Voyager builds a skill database for low-level control, while DEPS uses a structured four-step planning process (Describe, Explain, Plan, Select). However, these methods do not truly learn: they store successful plans or skills and use primitive composition (retrieval, recombination) for future tasks. DEPS cannot analyze why plans succeed or fail, and Voyager’s skill library ignores causal dependencies. Both treat experience as static data, limiting generalization and adaptation.
@@
  • A cognitive-inspired formulation of artificial mental models to enable natural-language memory accumulation and transfer learning,
  • MINDSTORES, a novel open-world LLM planner leveraging the above formulation to develop lifelong learning embodied agents, and
@@
In the remainder of this paper, we detail the theoretical foundations of mental models in Section  2, present the MINDSTORES architecture in Section  3, and validate its performance through experiments in Sections  4 and  5. Our findings underscore the critical role of memory-informed reasoning in developing lifelong learning agents for open-world environments.
@@ 2.1 Open-World Planning for Embodied Agents
Planning for embodied agents in open-world environments presents unique challenges due to the unbounded action space, long-horizon dependencies, and complex environmental dynamics. In environments like Minecraft, agents must reason about sequences of actions that may span dozens of steps, where early mistakes can render entire trajectories infeasible (Fan et al., 2022). Traditional planning approaches that rely on explicit state representations and value functions struggle in such domains due to the combinatorial explosion of possible states and actions. The key challenges in open-world planning stem from two main factors. First, the need for accurate multi-step reasoning due to long-term dependencies between actions presents a significant hurdle. Second, the requirement to consider the agent’s current state and capabilities when ordering parallel sub-goals within a plan poses additional complexity. Consider the example of crafting a diamond pickaxe in Minecraft: the process requires first obtaining wood, then crafting planks and sticks, mining stone with a wooden pickaxe, crafting a stone pickaxe, mining iron ore, smelting iron ingots, and finally crafting the iron pickaxe – a sequence that can easily fail if any intermediate step is incorrectly executed or ordered.
@@ 2.2 Zero-Shot LLM Planning with DEPS
Recent work has shown that large language models can serve as effective zero-shot planners for embodied agents through their ability to decompose high-level tasks into sequences of executable actions (Huang et al., 2022a). The DEPS (Describe, Explain, Plan and Select) framework leverages this capability through an iterative planning process that combines several key components (Wang et al., 2024c). The framework utilizes a descriptor that summarizes the current state and execution outcomes, an explainer that analyzes plan failures and suggests corrections, a planner that generates and refines action sequences, and a selector that ranks parallel candidate sub-goals based on estimated completion steps. The key innovation of DEPS is its ability to improve plans through verbal feedback and explanation. When a plan fails, the descriptor summarizes the failure state, the explainer analyzes what went wrong, and the planner incorporates this feedback to generate an improved plan. This creates a form of zero-shot learning through natural language interaction. However, DEPS and similar approaches maintain no persistent memory across episodes. Each new planning attempt starts fresh, unable to leverage insights gained from previous successes and failures in similar situations. This limitation motivates our work on experience-augmented planning.
@@ 2.3 Mental Models
Mental models are cognitive representations of how systems and environments work, enabling humans to understand, predict, and interact with the world around them. Originally proposed by Craik (1952), mental models theory suggests that people construct small-scale internal models of reality that they use to reason, anticipate events, and guide behavior. These models are built through experience and observation, continuously updated as new information becomes available, and help reduce cognitive load by providing ready-made frameworks for understanding novel situations. A key insight from psychological research on mental models is their role in transfer learning and generalization (Canini et al., ). When faced with new scenarios, humans naturally draw upon their existing mental models to make informed decisions, even in previously unseen contexts. This ability to leverage past experiences through abstract representations is particularly relevant for embodied agents operating in open-world environments, where they must constantly adapt to novel situations while maintaining coherent, generalizable knowledge about environmental dynamics.
@@ 3.3 Experience-Guided Planning
$\mathcal{N}_{k}(\mathcal{D},s_{t},t_{t})=\text{top-k}_{(s,t,p,o)\in\mathcal{D}}[\sum_{x\in\{s,t\}}\lambda_{x}\text{sim}(x,x_{t})]$ (2)
@@
$\mathcal{N}_{k}(s_{t},t_{t},p_{t})=\text{top-k}_{(s,t,p,o)\in\mathcal{D}}[\sum_{x\in\{s,t,p\}}\lambda_{x}\text{sim}(x,x_{t})]$ (3)
@@
0:  State $s_{t}$, Task $t_{t}$, Database $\mathcal{D}$, LLM $M$, $k$ neighbors
0:  Plan $p_{t}$
1:  $\mathcal{N}_{k}\leftarrow\texttt{retrieve\_top\_k}(\mathcal{D},s_{t},t_{t},k)$
2:  $\text{insights}\leftarrow M.\texttt{analyze\_experiences}(\mathcal{N}_{k})$
3:  $p_{t}\leftarrow M.\texttt{generate\_plan}(s_{t},t_{t},\mathcal{N}_{k},\texttt{insights})$
4:  while true do
5:     similar_plans $\leftarrow$
@@ 3.4 Design Justification
Our approach incorporates several carefully considered design elements that work together to create an effective planning system. The use of natural language experiences, rather than vectors or symbolic representations, leverages the LLM’s ability to perform flexible reasoning over arbitrary descriptions. The semantic retrieval system employs dense embeddings to enable efficient similarity search while capturing semantic relationships between experiences beyond exact matches. The two-stage retrieval process first retrieves experiences based on state/task similarity to inform plan generation, then retrieves similar plans to predict outcomes, allowing the planner to both learn from past experiences and validate new plans. Finally, the iterative refinement capability enables the planner to revise plans based on predicted outcomes before execution, reducing the cost of actual failures. This design maintains the benefits of zero-shot LLM planning while enabling continual learning through natural experience.
@@ 4.1 Experimental Setup
We evaluate our experience-augmented planning approach in MineDojo using 8 tiers of task complexity complexity (MT1-MT8) (Fan et al., 2022). The observation space includes RGB view, GPS coordinates, and inventory state, with 42 discrete actions mapped from MineDojo’s action space (Fan et al., 2022). All experiments utilize the behavior cloning controller trained on human demonstrations, following similar methodology to DEPS and Voyager. Due to software version constraints, our implementation of the controller achieves lower baseline performance than the original DEPS controller. Therefore, we use our implementation of DEPS without the experience database as the primary baseline for fair comparison. Each task is evaluated over 30 trials with randomized initial states and a fixed random seed of 42.
@@ 4.2 Evaluation Tasks
  • Basic (MT1-MT2): Fundamental tasks (wood/stone tools, basic blocks)
  • Intermediate (MT3-MT5): Progressive Tasks (food, mining, armor crafting)
  • Advanced (MT6-MT8): Complex tasks (iron tools, minecart, diamond)
@@ 4.5 Metrics
We measure:
@@
  • Complexity Scaling: Performance vs task complexity tiers
  • Retrieval Impact: Success rate vs context size (k)
@@ 5.1 Performance Metrics
MT1MT2MT3MT4MT5MT6MT7MT80020202020404040406060606080808080100100100100Success Rate (%)MINDSTORESDEPS
@@
As we analyze Figure 3 in comparison to our version of DEPS, we see all-around improvement with the experience database addition.
@@ 5.2 Learning Efficiency Analysis
MINDSTORES demonstrates superior learning efficiency, particularly for complex tasks. For basic tasks like mining wood and cobblestone, all systems perform comparably (9 to 42 iterations) (Figure 4). However, as complexity increases, MINDSTORES requires fewer iterations (54 to 276) compared to Voyager and Reflexion, which show exponential increases in required iterations.
Mine WoodMine CobblestoneMine CoalFurnaceMake Stone SwordMine Iron00100100100100200200200200300300300300400400400400500500500500Novel Learning IterationsMINDSTORESVoyagerReflexion
Figure 4: Performance comparison between MINDSTORES, Voyager, and Reflexion across different Minecraft tasks. Values capped at 500 iterations (shown for Reflexion in later tasks). MINDSTORES demonstrates superior efficiency in complex tasks. (Note: Novel Learning Iterations refer to the amount of unseen additions to the experience database)
@@ 5.3 Scalability with Task Complexity
Performance divergence becomes pronounced with increasing task complexity. MINDSTORES maintains efficient novel learning iterations for tasks like crafting a stone sword and mining iron, while Voyager and Reflexion require significantly more iterations, even reaching the max range (500+) for a relatively simple Mine Iron task (Figure 4).
[…]
Figure 5: Success rates vs retrieval context size $k$ for different tasks
5.4 Performance Scaling with $k$
Success rates improve significantly as $k$ increases from 1 to 10 but show diminishing returns at $k=20$. The impact of $k$ varies by task complexity (Figure 5): for simple tasks such as Torch crafting, success rates show steady improvement from 3.3% to 23.3% up to $k=10$. Medium-complexity tasks like Iron Boots exhibit more gradual improvement, with success rates rising from 10% to 33.3%. For complex tasks such as Iron Pickaxe and Minecart crafting, feasibility is only achieved with larger $k$ values, highlighting the dependency on increased computational resources. However, end-game tasks like Diamond crafting remain unachievable regardless of the value of $k$, underscoring the inherent limitations of the system in handling highly complex objectives.
These results align with our expected theorized improvements from the Voyager and DEPS architectures, highlighting the exponential impact of task complexity on completion times in open-world environments. While basic operations are handled reliably, managing complex, multi-step tasks remains a challenge.
@@ 5.5 Continuous Experience Building Analysis
Within Figure 6, we present findings which shows an experiment in which we do not reset the experience database after each task is queried, but instead, we keep the experience database building such that we see the effect of a “global” experience database over a multitude of tasks. We observe exceptional results, with the entire process of completing the Minecart task taking only 9112 steps including the previous 9 tasks. This fully outperforms the allotted 6000 steps needed to complete the task in a new environment, showing that we only required  200 new steps. New task completion steps (Step differential between each following task) decrease non-linearly even as complexity grows:
@@
The system maintains a 100% success rate across all tasks, indicating robust skill transfer and knowledge utilization from the growing experience database, which expands from 26 entries for wooden door to 355 entries for minecart (Figure 6 and Appendix Table 4).
[…]
Figure 6: Steps required for task completion with continuous building of experience database. See Appendix Table 4 for corresponding tasks.
@@ 6.1 Embodied Planning & Classical Methods
Early approaches used hierarchical reinforcement learning (Sutton et al., 1999) and symbolic planning (Kaelbling & Lozano-Perez, 2011) but struggled with scalability in open-world domains like Minecraft. Hybrid methods like PDDLStream (Garrett et al., 2020) combined symbolic planning with procedural samplers, while DreamerV3 (Hafner et al., 2024) employed latent world models. However, these methods depend on rigid priors, lack causal reasoning, and fail to recover from irreversible errors. Reinforcement learning frameworks (e.g., DQN (Mnih et al., 2015), PPO (Schulman et al., 2017)) and LLM-RL hybrids like Eureka (Ma et al., 2023) also falter in dynamic, long-horizon tasks due to static reward mechanisms and error propagation.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated enhanced capabilities in reasoning (Plaat et al., 2024; Huang & Chang, 2023), planning (Sel et al., 2025), and decision-making (Huang et al., 2024) through methods that strengthen analytical depth. Among the numerous domains of active innovation, the success of AI agents serves as a critical benchmark for assessing our progress toward generally capable artificial intelligence (Brown et al., 2020).
@@
A promising paradigm for such agents leverages LLMs as high-level planners (Jeurissen et al., 2024): the LLM decomposes abstract goals into step-by-step plans (e.g., “mine wood $\rightarrow$ craft tools $\rightarrow$ smelt iron”), while a low-level controller translates these plans into environment-specific actions (e.g., movement, object interaction). This “brain and body” architecture capitalizes on the LLM’s capacity for structured reasoning while grounding its outputs in the dynamics of the physical world—a critical capability for real-world applications like robotic manipulation (Shentu et al., 2024; Bhat et al., 2024; Wang et al., 2024b), autonomous navigation (Zawalski et al., 2024), and adaptive disaster response.
InputReview ExperiencePlan GenerationPredict Outcome ExecuteRecordFailure ManagementNext IterationTask: Mine IronReview: Previous experiences for mining needs a pickaxe!Plan: Mine Refer to caption, craft Refer to caption, mine Refer to caption, craft Refer to caption, mine Refer to captionPredict: Probable failure, might die from hunger while finding iron ore.Review: Find food before searching for iron ore.Plan: Mine Refer to caption, craft Refer to caption, mine Refer to caption, craft Refer to caption, craft Refer to caption, hunt Refer to caption, mine Refer to captionPredict: Probable success!
@@
Figure 1: Overview of the MINDSTORES planning architecture. The left shows the iterative experiential learning pipeline leveraging the experience database. Database-related methods are in orange, planning steps are in green, and Minecraft steps are in red. The right shows an example applying this pipeline to an example task in Minecraft.
@@
  • A cognitive-inspired formulation of artificial mental models to enable natural-language memory accumulation and transfer learning.
  • MINDSTORES, a novel open-world LLM planner leveraging the above formulation to develop lifelong learning embodied agents.
@@
In the remainder of this paper, we detail the theoretical foundations of mental models in Section 2, present the MINDSTORES architecture in Section 3, and validate its performance through experiments in Sections 4 and 5. Our findings underscore the critical role of memory-informed reasoning in developing lifelong learning agents for open-world environments.
@@ 2.1 Open-World Planning for Embodied Agents
Planning for embodied agents in open-world environments presents unique challenges due to the unbounded action space, long-horizon dependencies, and complex environmental dynamics. In environments like Minecraft, agents must reason about sequences of actions that may span dozens of steps, where early mistakes can render entire trajectories infeasible (Fan et al., 2022). Traditional planning approaches that rely on explicit state representations and value functions struggle in such domains due to the combinatorial explosion of possible states and actions.
The key challenges in open-world planning stem from two main factors. First, the need for accurate multi-step reasoning due to long-term dependencies between actions presents a significant hurdle. Second, the requirement to consider the agent’s current state and capabilities when ordering parallel sub-goals within a plan poses additional complexity. Consider the example of crafting a diamond pickaxe in Minecraft: the process requires first obtaining wood, then crafting planks and sticks, mining stone with a wooden pickaxe, crafting a stone pickaxe, mining iron ore, smelting iron ingots, and finally crafting the iron pickaxe—a sequence that can easily fail if any intermediate step is incorrectly executed or ordered.
@@ 2.2 Zero-Shot LLM Planning with DEPS
Recent work has shown that large language models can serve as effective zero-shot planners for embodied agents through their ability to decompose high-level tasks into sequences of executable actions (Huang et al., 2022a). The DEPS (Describe, Explain, Plan and Select) framework leverages this capability through an iterative planning process that combines several key components (Wang et al., 2024c). The framework utilizes a descriptor that summarizes the current state and execution outcomes, an explainer that analyzes plan failures and suggests corrections, a planner that generates and refines action sequences, and a selector that ranks parallel candidate sub-goals based on estimated completion steps.
The key innovation of DEPS is its ability to improve plans through verbal feedback and explanation. When a plan fails, the descriptor summarizes the failure state, the explainer analyzes what went wrong, and the planner incorporates this feedback to generate an improved plan. This creates a form of zero-shot learning through natural language interaction.
However, DEPS and similar approaches maintain no persistent memory across episodes. Each new planning attempt starts fresh, unable to leverage insights gained from previous successes and failures in similar situations. This limitation motivates our work on experience-augmented planning.
@@ 2.3 Mental Models
Mental models are cognitive representations of how systems and environments work, enabling humans to understand, predict, and interact with the world around them. Originally proposed by Craik (1952), mental models theory suggests that people construct small-scale internal models of reality that they use to reason, anticipate events, and guide behavior. These models are built through experience and observation, continuously updated as new information becomes available, and help reduce cognitive load by providing ready-made frameworks for understanding novel situations.
A key insight from psychological research on mental models is their role in transfer learning and generalization (Canini et al., 2010). When faced with new scenarios, humans naturally draw upon their existing mental models to make informed decisions, even in previously unseen contexts. This ability to leverage past experiences through abstract representations is particularly relevant for embodied agents operating in open-world environments, where they must constantly adapt to novel situations while maintaining coherent, generalizable knowledge about environmental dynamics.
@@ 3.3 Experience-Guided Planning
$\mathcal{N}_{k}(\mathcal{D},s_{t},t_{t})=\text{top-k}_{(s,t,p,o)\in\mathcal{D}}\left[\sum_{x\in\{s,t\}}\lambda_{x}\text{sim}(x,x_{t})\right]$ (2)
@@
$\mathcal{N}_{k}(s_{t},t_{t},p_{t})=\text{top-k}_{(s,t,p,o)\in\mathcal{D}}\left[\sum_{x\in\{s,t,p\}}\lambda_{x}\text{sim}(x,x_{t})\right]$ (3)
@@
1:State $s_{t}$, Task $t_{t}$, Database $\mathcal{D}$, LLM $M$, $k$ neighbors
2:Plan $p_{t}$
3:$\mathcal{N}_{k}\leftarrow\texttt{retrieve\_top\_k}(\mathcal{D},s_{t},t_{t},k)$
4:$\text{insights}\leftarrow M.\texttt{analyze\_experiences}(\mathcal{N}_{k})$
5:$p_{t}\leftarrow M.\texttt{generate\_plan}(s_{t},t_{t},\mathcal{N}_{k},\texttt{insights})$
6:while true do
@@ 3.4 Design Justification
Our approach incorporates several carefully considered design elements that work together to create an effective planning system. The use of natural language experiences, rather than vectors or symbolic representations, leverages the LLM’s ability to perform flexible reasoning over arbitrary descriptions. The semantic retrieval system employs dense embeddings to enable efficient similarity search while capturing semantic relationships between experiences beyond exact matches. The two-stage retrieval process first retrieves experiences based on state/task similarity to inform plan generation, then retrieves similar plans to predict outcomes, allowing the planner to both learn from past experiences and validate new plans. Finally, the iterative refinement capability enables the planner to revise plans based on predicted outcomes before execution, reducing the cost of actual failures.
This design maintains the benefits of zero-shot LLM planning while enabling continual learning through natural experience.
@@ 4.1 Experimental Setup
We evaluate our experience-augmented planning approach in MineDojo using 8 tiers of task complexity (MT1–MT8) (Fan et al., 2022). The observation space includes RGB view, GPS coordinates, and inventory state, with 42 discrete actions mapped from MineDojo’s action space (Fan et al., 2022). All experiments utilize the behavior cloning controller trained on human demonstrations, following similar methodology to DEPS and Voyager. Due to software version constraints, our implementation of the controller achieves lower baseline performance than the original DEPS controller. Therefore, we use our implementation of DEPS without the experience database as the primary baseline for fair comparison. Each task is evaluated over 30 trials with randomized initial states and a fixed random seed of 42.
@@
For the complete agent algorithm and associated LLM prompts, see Appendix A, and for detailed implementation aspects including environment integration and neural component configurations, see Appendix B.
@@ 4.2 Evaluation Tasks
  • Basic (MT1–MT2): Fundamental tasks (wood/stone tools, basic blocks)
  • Intermediate (MT3–MT5): Progressive tasks (food, mining, armor crafting)
  • Advanced (MT6–MT8): Complex tasks (iron tools, minecart, diamond)
@@
For additional task details and performance statistics, see Appendix C.
@@ 4.3 Baselines
We compare MINDSTORES to the following existing state-of-the-art approaches:
@@ 4.4 Ablations
To analyze the function of each individual component of the MINDSTORES framework, we perform the following ablations:
@@ 4.5 Metrics
To quantify each method’s performance in open-world planning in the Minecraft environment, we measure:
@@
  • Complexity Scaling: Performance vs. task complexity tiers
  • Retrieval Impact: Success rate vs. context size ($k$)
@@ 5.1 Performance Metrics
1234567800101010102020202030303030404040405050505060606060707070708080808090909090100100100100Task Difficulty (MT)Success Rate (%)MINDSTORESDEPS
@@
WoodCobblestoneCoalFurnaceSwordIron0050505050100100100100150150150150200200200200250250250250300300300300350350350350400400400400450450450450500500500500Novel Learning IterationsMINDSTORESVoyagerReflexion
Figure 4: Novel learning iterations across different Minecraft tasks. MINDSTORES demonstrates superior efficiency in complex tasks. (Note: Iteration counts for Reflexion are capped at 500 in later tasks.)
As we analyze Figure 4 in comparison to our version of DEPS, we see an all-around improvement with the addition of the experience database.
@@ 5.2 Learning Efficiency Analysis
MINDSTORES demonstrates superior learning efficiency, particularly for complex tasks. For basic tasks like mining wood and cobblestone, all systems perform comparably (9 to 42 iterations) (see Figure 4). However, as complexity increases, MINDSTORES requires fewer iterations (54 to 276) compared to Voyager and Reflexion, which show exponential increases in required iterations.
[…]
Figure 5: Success rates vs. retrieval context size $k$ for different tasks. Simple tasks improve steadily with $k$, while more complex tasks require larger $k$ values. Advanced tasks remain unachievable regardless of $k$.
[…]
Figure 6: Steps required for task completion with continuous building of the experience database. (See Appendix Table 1 for corresponding tasks.)
@@ 5.3 Scalability with Task Complexity
Performance divergence becomes pronounced with increasing task complexity. MINDSTORES maintains efficient novel learning iterations for tasks like crafting a stone sword and mining iron, while Voyager and Reflexion require significantly more iterations, even reaching the max range (500+) for a relatively simple Mine Iron task (see Figure 4).
@@ 5.4 Continuous Experience Building Analysis
Figure 6 shows an experiment in which the experience database is not reset between tasks but is built continuously across multiple tasks. We observe that the entire process of completing the Minecart task takes only 9112 steps including the previous 9 tasks, compared to the 6000 steps needed in a fresh environment. This indicates that only approximately 200 new steps were required. The number of new task completion steps decreases non-linearly even as task complexity grows:
@@
The system maintains a 100% success rate across all tasks, indicating robust skill transfer and knowledge utilization from the growing experience database, which expands from 26 entries for Wooden Door to 355 entries for Minecart (see Figure 6 and Appendix Table 1).
@@ 6.1 Embodied Planning & Classical Methods
Early approaches used hierarchical reinforcement learning (Sutton et al., 1999) and symbolic planning (Kaelbling & Lozano-Perez, 2011) but struggled with scalability in open-world domains like Minecraft. Hybrid methods like PDDLStream (Garrett et al., 2020) combined symbolic planning with procedural samplers, while DreamerV3 Hafner et al. (2024) employed latent world models. However, these methods depend on rigid priors, lack causal reasoning, and fail to recover from irreversible errors. Reinforcement learning frameworks (e.g., DQN (Mnih et al., 2015), PPO (Schulman et al., 2017)) and LLM-RL hybrids like Eureka (Ma et al., 2023) also falter in dynamic, long-horizon tasks due to static reward mechanisms and error propagation.